OpenStudy (anonymous):

X(y-z) if x = -2,y = 3, and z = -1

OpenStudy (anonymous):

\(x(y-z)\) \(=(-2)(3-(-1))\\ =-2(3+1)\\ =-2(4)\\ =-8\)

OpenStudy (anonymous):

Thank u true5

OpenStudy (anonymous):

Two ways to solve 10(4 + 1/2)

OpenStudy (anonymous):

Add the terms together in parenthesis: 4+1/2 = 4.5 Then multiply ... 10(4.5) = 45

OpenStudy (anonymous):

Or use distributivity: 10(4) + 10(1/2)= 40 + 5 = 45